# SEKE: Specialised Experts for Keyword Extraction

### v1

@@ 5 Results and Analysis
In Table 2, we present the results achieved by our approach and compare them to several baselines (see Section 2.2 for details). TF-IDF (we list results from Martinc et al. (2022)), TextRank (we list results from Martinc et al. (2022)), ChatGPT (we report results for the gpt-3.5-turbo version tested in Luo et al. (2023)) and ChatGLM2-6B (we list results from Luo et al. (2023)) algorithms are unsupervised and do not require any training. For the supervised baselines, CopyRNN, CatSeqD, GPT-2, GPT-2+BiLSTM-CRF, and TNT-KID, we list results reported by Martinc et al. (2022), for Diff-KPE, we list results reported by Luo et al. (2023) and for JointKPE, we list results from Sun et al. (2021). Besides testing our custom sequence labelling head (containing MoE, RNN, or both), we also report results for the unmodified pre-trained DeBERTa model with a standard feed-forward token classification head. For it, we apply the same fine-tuning regime as for our experiments with a custom classification head.
@@
In the validation set training scenario, by adding the MoE on top of the DeBERTa model (see configuration DeBERTa MoE), we improve the performance of the vanilla model (see configuration DeBERTa) with a feed-forward token classification head on four out of six datasets according to all evaluation criteria. The improvement is not observed for the two smallest datasets, Krapivin and Inspec. On the other hand, adding the additional RNN head to the vanilla DeBERTa model improves the performance of the base model on all datasets (see configuration DeBERTa RNN) and according to almost all criteria (the only exception being the F1@1 criterion on the openKP dataset). The results therefore indicate a synergy between the MoE and RNN layers, since the model with both MoE and RNN layers manages to substantially improve the base model and also outperforms the DeBERTa MoE and DeBERTa RNN models on four out of six datasets according to all criteria.
@@
In the full training scenario, the DeBERTa MoE model outperforms the vanilla DeBERTa on three out of six datasets according to all evaluation criteria. Interestingly, the DeBERTa MoE model fails on all computer science datasets, which suggests that expert specialisation might be harder on this domain. On the other hand, the DeBERTa MoE RNN model outperforms the vanilla one on all datasets and according to almost all criteria, the only exception again being the F1@1 criterion on the openKP dataset.
@@
In this scenario we also manage to substantially improve the SOTA on the Inspec, KPTimes, and JPTimes datasets according to both criteria. While we outperform the best-performing model JoinKPE in F1@10 on the KP20k and the Krapivin datasets, our best-performing configuration (DeBERTa MoE RNN) on the other hand achieves uncompetitive performance in terms of F1@5 on these two datasets. On the openKP, our best approach (DeBERTa MoE RNN) fails to beat JointKPE and Diff-KPE approaches according to the F1@1 criterion. The performance comparison is nevertheless favourable according to the other two criteria. Finally, the results also show that the unsupervised models in general achieve much worse performance than the supervised models.
@@ 5.1 Examining the Specialisation of Experts
Table 3: Results of the Cramer’s V statistic for correlation between the most weighted expert for a specific token and the specialisation categories examined, across Validation set training and Full training regimes.
@@
In the validation set training scenario, a strong correlation between specific words and specific experts is observed on all datasets. This is not surprising, since MoE works on the token level. On the other hand, the correlation between experts and labels (I, O, and B) tends to be moderate at best (i.e., between 0.1 and 0.3) on most datasets. One can also observe a moderate (and on some datasets strong, i.e., more than 0.5) correlation between specific experts and POS tags, suggesting that specialisation is to a large extent influenced by the syntax of the text.
@@
By looking at the other three types of correlation, namely the correlation between experts and NEs, between experts and stopword tokens, and between experts and punctuation tokens, one can see a very distinct difference between the computer science and other datasets. While there is a strong tendency for experts to specialize for stopword and punctuation tokens on all the computer science datasets (Inspec, Krapivin, and KP20k), this tendency is much weaker on other (news and web page) datasets, where one can observe almost no correlation. On the contrary, here one can observe a specialisation of experts for NEs, which is almost non-existent on the computer science datasets. A correlation is almost the same if we binarize the NE sequence (i.e., to NE/non-NE tokens) or if we look for correlations between experts and 23 labels returned by the SpaCy NE recognition pipeline. The specialisation for NEs can be perhaps explained by the fact that news datasets contain much more NEs than scientific papers and that many of them are in fact labelled as keywords. The same is true for the OpenKP dataset that includes web pages sampled from the index of Bing search engine content-oriented pages, i.e., news pages, multi-media pages from video sites, or indexing pages. The strong correlation with simple syntactic patterns on the computer science datasets can be on the other hand explained by the lack of obvious semantic clues, such as NEs, in these datasets.
@@
When comparing the specialisation in the validation set training scenario and the full training scenario, one can see that correlation to labels, words, NEs, and binary NEs is mostly the same. On the other hand, one can see a substantial decrease in correlation to POS tags, stopwords and punctuation on the computer science datasets, when the models are trained on more data. The assumption is that this is the consequence of expert specialisation becoming more complex when the amount of data allows it.
@@ 6 Conclusion
In summary, we investigated the potential of MoE for keyword extraction tasks. We introduced SEKE, a novel architecture that combines MoE and RNN built upon the DeBERTa model for English. This approach achieved superior performance compared to existing SOTA models on standard datasets for keyword extraction. Further analysis confirmed the benefits of MoE layers during fine-tuning, especially with adequate data. The study revealed that adding recurrent layers can help mitigate the data limitations, showing the feasibility of the approach even on small training datasets with less than 2,000 documents. Nevertheless, the specialization analysis shows that the size of the training data still influences the specialisation and that more data allows for more complex specialisation patterns to develop. In the future, we will try to pinpoint the exact dataset size thresholds, down to which the proposed approach is still feasible.
@@
The analysis also revealed that specialisation is domain-specific. Current results indicate that in text with higher complexity (i.e., scientific papers) experts to a larger extent rely on syntactical patterns than in the less complex text, where experts to a larger extent specialize for processing of semantic clues. This hypothesis will be further tested in the future by applying the approach on new domains and languages other than English. This expansion will also further test the generalisability of the proposed approach, reducing the language and domain limitations of the current study.
Another limitations is the usage of an BiLSTM encoder in the custom token classification head, which makes the inference and fine-tuning slightly slower and more computationally demanding than in the vanilla DeBERTa model. While the difference in time was negligible in all our experiments, it might have a visible effect, if the model is applied on a very large dataset. We should also point out the reliance on SpaCY in the specialisation analysis, which is a general domain NLP tool, rarely employed on specialized texts, such as scientific articles. Its usage might have affected the reliability of the analysis’ correlation results by extracting fewer or incorrect NEs.
@@
Finally, we can observe that despite the impressive improvements in the development of foundational LLMs that offer impressive zero-shot capabilities, keyword extraction task can still not be sufficiently solved by employing these models, which lag substantially behind most supervised approaches on most datasets. The biggest advantage of supervised approaches is their ability to adapt to the specifics of the syntax, semantics, content, genre and keyword tagging regime of the specific corpus Meng et al. (2017). Since this cannot be done in a zero-shot setting, in the future we will try to adapt LLMs to the specific corpora and keyword extraction regimes with fine-tuning and in-context learning.
7 Acknowledgements
@@
Appendix A Appendix: Examples of keyword identification

### v2

@@ 5 Results and Analysis
In Table 2, we present the results of our approach and compare them to several baselines (see Section 2.2 for details). TF-IDF, TextRank, RaKUN and YAKE algorithms are unsupervised and do not require any training. Same applies to the LLM-based approaches, namely ChatGPT [^3] We report results for the gpt-3.5-turbo version tested in Luo et al. (2024)., ChatGLM2-6B [^4] We list results from Luo et al. (2024)., GPT4o-mini (Achiam et al., 2023) and Gemini-1.5-flash-8B (Team et al., 2023)555see Section D for details about the LLM experiments..
For the supervised baselines, CopyRNN, CatSeqD, GPT-2, GPT-2+BiLSTM-CRF, and TNT-KID, we list results reported by Martinc et al. (2022), for Diff-KPE, we list results reported by Luo et al. (2024), for JointKPE, we list results from Sun et al. (2021), and for HybridMatch, we list results from Song et al. (2024). Note that GPT-2, GPT-2 + BiLSTM-CRF and TNT-KID are sequence labelling approaches, same as ours. Since they return a non-predetermined number of keywords (i.e., the approaches learn during training how many keywords they should return), they require adaptation to each specific keyword labelling regime for optimal performance and were therefore fine-tuned on the validation datasets. In contrast, JointKPE, DIffKPE nad HybridMatch baselines return a predetermined number of best ranked candidates (3, 5 or 10, depending on the evaluation criteria) and do not require validation set fine-tuning for adaptation. Same applies to the two generation baselines, CopyRNN and CatSeqD, which tend to generate a large set of predicted keywords, from which a fixed number of top predicted phrases (3, 5 or 10) is taken to compare with the ground truth. Finally, besides testing our custom sequence labelling heads (containing MoE, BiLSTM, or both), we also report results for the unmodified pre-trained DeBERTa model with a standard feed-forward token classification head.
@@
In the validation set training scenario, by adding the MoE on top of the DeBERTa model (see configuration DeBERTa MoE), we improve the performance of the vanilla model (see configuration DeBERTa) with a feed-forward token classification head on four out of six datasets according to all evaluation criteria. The improvement is not observed for the two smallest datasets, Krapivin and Inspec. On the other hand, adding the additional BiLSTM head to the vanilla DeBERTa model improves the performance of the base model on all datasets (see configuration DeBERTa BiLSTM) and according to almost all criteria (the only exception being the F1@1 criterion on the openKP dataset). The results therefore indicate a synergy between the MoE and BiLSTM layers, since the model with both MoE and BiLSTM layers manages to substantially improve the base model and also outperforms the DeBERTa MoE and DeBERTa BiLSTM models on four out of six datasets according to all criteria.
@@
In the full training scenario, the DeBERTa MoE model outperforms the vanilla DeBERTa on three out of six datasets according to all evaluation criteria. Interestingly, the DeBERTa MoE model fails on all computer science datasets, which suggests that expert specialisation might be harder on this domain. On the other hand, the DeBERTa MoE BiLSTM model outperforms the vanilla one on all but one dataset and criterion, the F1@1 criterion on the openKP dataset.
@@
In this scenario, we also manage to substantially improve the SOTA on the Inspec, KPTimes, and JPTimes datasets according to both criteria. While we outperform the best-performing model JoinKPE in F1@10 on the KP20k and the Krapivin datasets, our best-performing configuration (DeBERTa MoE BiLSTM) on the other hand achieves uncompetitive performance in terms of F1@5 on these two datasets. On the openKP, our best approach (DeBERTa MoE BiLSTM) fails to beat several approaches (among others, also the two LLM-based baselines, GPT4o-mini and Gemini-1.5-flash-8B) according to the F1@1 criterion. The performance comparison is nevertheless favourable according to the other two criteria.
The results also show that the unsupervised models in general achieve much worse performance than the supervised models. While the LLM-based approaches outperform other unsupervised baselines, their performance lags behind supervised approaches. Out of the four tested LLMs, the two proprietary models, GPT4o-mini and Gemini-1.5-flash-8B, perform the best, with Gemini-1.5-flash-8B being the best model on most datasets.
@@ 5.1 Examining the Specialisation of Experts
Table 3: Results of the Cramer’s V statistic for correlation between the most weighted expert for a specific token and the specialisation categories examined, across Validation set training and Full training regimes. Strongest, second strongest, and third strongest correlations per dataset and regime are marked.
@@
In the validation set training scenario, a strong correlation between specific words and specific experts is observed on all datasets. This is not surprising, since MoE works on the token level. In contrast, the correlation between experts and labels (I, O, and B) tends to be moderate at best (i.e., between 0.1 and 0.3) on most datasets. One can also observe a moderate (and on some datasets strong, i.e., more than 0.5) correlation between specific experts and POS tags, suggesting that specialisation is somewhat influenced by the syntax of the text.
@@
By looking at the other three types of correlation, namely the correlation between experts and NEs, between experts and stopword tokens, and between experts and punctuation tokens, one can see a very distinct difference between the computer science and other datasets. While there is a strong tendency for experts to specialise for stopword and punctuation tokens on all the computer science datasets (Inspec, Krapivin, and KP20k), this tendency is much weaker on other (news and web page) datasets, where one can observe almost no correlation. Here one can observe a specialisation of experts for NEs, which is almost non-existent on the computer science datasets. A correlation is almost the same if we binarize the NE sequence (i.e., to NE/non-NE tokens) or if we look for correlations between experts and 23 labels returned by the SpaCy NE recognition pipeline. The specialisation for NEs can be perhaps explained by the fact that news and web datasets contain much more NEs than scientific papers and that many of them are in fact labelled as keywords. The strong correlation with simple syntactic patterns on the computer science datasets can be on the other hand explained by the lack of obvious semantic clues, such as NEs, in these datasets.
@@
When comparing the specialisation in the validation set training scenario and the full training scenario, one can see that correlation to labels, words, NEs, and binary NEs is mostly the same. On the other hand, one can see a decrease in correlation to POS tags, stopwords and punctuation on the computer science datasets, when the models are trained on more data. The assumption is that this is the consequence of expert specialisation becoming more complex when the amount of data allows it.
5.2 Error Analysis of MoE Performance by Keyword Length
Dataset/Keyword Length 1 2 3 4 5 6 7 8 9 10
Inspec 0.0018 (n=291)
-0.0042
(n=472)
0.0136 (n=396) 0.0303 (n=209) 0.0353 (n=85) 0.2610 (n=23) 0.1004 (n=10) 0.2381 (n=3)
0.0000
(n=1)
0.0000
(n=1)
JPTimes 0.0175 (n=9241)
-0.0386
(n=6642)
-0.0362
(n=1576)
0.0036 (n=274) 0.0417 (n=48) 0.1333 (n=15) 0.2500 (n=4)
0.0000
(n=4)
0.0000
(n=1)
KP20k
-0.0196
(n=9479)
0.0035 (n=14905) 0.0202 (n=7943) 0.0675 (n=2040) 0.0527 (n=450) 0.0103 (n=97)
0.0000
(n=27)
0.0000
(n=6)
0.0000
(n=5)
KPTimes 0.0222 (n=7184) 0.0814 (n=5027) 0.1121 (n=1766) 0.1569 (n=465) 0.1192 (n=151) 0.0909 (n=88) 0.3824 (n=34) 0.5556 (n=9)
0.0000
(n=1)
0.0000
(n=2)
Krapivin 0.0002 (n=206) 0.0139 (n=381) 0.0273 (n=174) 0.0063 (n=50)
0.0000
(n=8)
0.0000
(n=6)
openKP 0.0180 (n=3125) 0.0311 (n=4118) 0.0211 (n=1956) 0.0159 (n=599) 0.0072 (n=139) 0.0385 (n=26)
0.0000
(n=5)
0.0000
(n=4)
0.0000
(n=1)
Table 4: Difference in F1@1 scores between DeBERTa MoE BiLSTM and DeBERTa MoE (DeBERTa MoE BiLSTM F1@1 - DeBERTa MoE F1@1) on keywords of different word length across different datasets.
To test the hypothesis that the MoE component struggles with longer keywords due to context slicing, we have conducted an additional error analysis. This ablation study compares the performance of DeBERTa MoE BiLSTM and DeBERTa MoE model on keywords of varying lengths.
Specifically, for each test dataset, we created subsets of documents containing keywords of a specific length, ranging from one to ten words. We then evaluated the performance in terms of F1@1 score for each model on these subsets, considering only the keywords of that specific length (we use F1@1 due to small set of keywords of specific length per document).
Table 4 presents the differences in F1@1 scores (calculated as DeBERTa MoE BiLSTM F1@1 - DeBERTa MoE F1@1). Positive values indicate that DeBERTa MoE BiLSTM outperformed DeBERTa MoE on keywords of specific length, while negative values indicate the opposite. The results show that the standard MoE model is indeed less effective at extracting long, multi-word keywords compared to the model with an BiLSTM encoder (i.e., the difference is generally bigger on longer keywords). This analysis supports the hypothesis that context slicing in the MoE architecture is a key deficiency.
@@ 6 Conclusion
In summary, we investigated the potential of MoE for keyword extraction tasks. We introduced SEKE, a novel architecture that combines MoE and BiLSTM built upon the DeBERTa model. This approach achieved superior performance compared to existing SOTA models on standard datasets for keyword extraction. The study revealed that adding BiLSTM can help mitigate the data limitations, showing the feasibility of the approach even on small training datasets. In the future, we will try to pinpoint the exact dataset size thresholds, down to which the proposed approach is still feasible. Additionally, we will expand the evaluation to more benchmarks to further strengthen the analysis.
In this work, our primary goal was to evaluate the general effectiveness of our proposed framework using a consistent and reasonable configuration across all experiments. To maintain a fair comparison and avoid overfitting to specific datasets, we used commonly adopted default values (e.g., 4 experts, top-k = 2) without extensive hyperparameter tuning. In the future, a more exhaustive hyperparameter search will be conducted, which could yield additional performance gains and further insights into the model behaviour.
@@
The analysis revealed that specialisation is domain-specific. Current results indicate that in text with higher complexity (i.e., scientific papers) experts to a larger extent rely on syntactical patterns than in the less complex text, where experts to a larger extent specialise for processing of semantic clues. This hypothesis will be further tested in the future by applying the approach in new domains and languages other than English.
@@
Finally, we can observe that despite the impressive improvements in the development of foundational LLMs, keyword extraction task can still not be sufficiently solved by employing these models, which lag behind supervised approaches on most datasets. Since the biggest advantage of supervised approaches is their ability to adapt to the specifics of the syntax, semantics, content, and keyword tagging regime of the specific corpus Meng et al. (2017), in the future we will adapt LLMs to the specific corpora and keyword extraction regimes with fine-tuning and in-context learning.
7 Limitations
First, the empirical evaluation was conducted exclusively on English-language datasets from three specific domains (computer science, news, and the web). Consequently, the generalizability of our findings to other languages and more diverse or specialized domains remains to be thoroughly tested.
Furthermore, the model’s internal mechanisms may be susceptible to certain weaknesses. One potential issue is an over-reliance on specific Part-of-Speech (POS) patterns for keyword identification. The model might learn heuristic shortcuts, associating keywords primarily with common grammatical structures (e.g., noun-noun compounds or adjective-noun phrases). This could limit its ability to identify valid but unconventional keywords and may lead to false positives. A related concern, specific to the MoE architecture, is the risk of the expert routing mechanism overfitting to superficial cues. The router might learn to dispatch tokens based on shallow features like capitalization or token position rather than deep semantic content, which would prevent true expert specialization and harm the model’s robustness on out-of-distribution data.
From a computational and environmental standpoint, the introduction of the BiLSTM encoder in the custom token classification head increases the model’s complexity. This makes both inference and fine-tuning slightly slower and more computationally demanding than the vanilla DeBERTa model. While the performance difference was negligible in our experiments, it could become a significant factor when scaling to much larger datasets. More broadly, the training of large models like DeBERTa is an energy-intensive process with an associated environmental cost, a factor that is critical to consider in the development of ever-larger architectures.
Methodologically, our specialization analysis relied on spaCy for Named Entity Recognition. As a general-domain NLP tool, its performance can be suboptimal on specialized texts like scientific articles. This reliance might have affected the reliability of the analysis by yielding fewer or less accurate named entities, potentially skewing the correlation results.
Finally, it is crucial to consider the broader implications and potential for misuse. A model effective at automated keyword extraction could be exploited for malicious purposes, such as "black-hat" Search Engine Optimization (SEO). Such a system could be used to generate spammy keywords to artificially boost the ranking of low-quality content, thereby degrading the integrity of information retrieval systems.
8 Ethics Statement
This work does not pose any ethical issues. All the data and tools used in this paper are publicly available under the MIT license. No private data or non-public information is used in this work.
Acknowledgments
@@
Appendix A Hyperparameters Used
We use the same hyperparameter setting across all datasets, namely, the following values were used during fine-tuning: learning rate of 2e-4, LoRA intrinsic rank of 16, LoRA alpha of 16, LoRA dropout of 0.1 sequence length of 256, 4 experts, and the top k value of 2.
Appendix B Model Size and Budget
The final model with MoE and BiLSTM layers, which uses the “deberta-v3-base” backbone (available at https://huggingface.co/microsoft/deberta-v3-base under the MIT license) has 274M parameters, out of them 45M are trainable during LoRA fine-tuning.
We used two NVIDIA A100 80GB GPUs in our experiments. The total computational budget for all experiments, i.e., validation set training and full training with five random seeds and four different experimental settings (feed-forward classification head, just MoE layer, just BiLSTM layer, both MoE and BiLSTM layers), was 1400 GPU hours.
Appendix C Examples of Keyword Identification
@@
Appendix D Prompt Used for LLM-based Keyword Extraction
For each document processed in our experiments, we applied the same prompt to ensure consistency in the extraction of relevant keywords. We condition the generation to return valid json prompts, so we can uniformly parse and compare the results. The prompt used for both the GPT4o-mini and Gemini-1.5-flash-8B models is provided below:
{internallinenumbers*} You are an assistant that specializes in text analysis. Your task is to extract the most relevant keywords from the following document. {internallinenumbers*} Extract only keywords present in the document. Please analyze the text and provide a comma-separated list of keywords ordered from most important to least important. Return a maximum of 10 keywords. Document: \"\"\"{document}\"\"\" {internallinenumbers*} Return only the list of keywords, formatted as a json list (i.e. with square brackets [ and ]).
Appendix E Performance of our approaches trained solely on training sets
To make a comparison between our approach and other baselines (i.e., especially non-sequence labelling supervised baselines, which were not fine-tuned on the validation data) more comprehensive, Table 5 presents results when training of our approaches is done solely on KP20k, KPTimes, and OpenKP, with no validation data. Note that in this training setting our approaches outperform all baselines in 7 out of 13 cases. On the other hand, a large drop in performance can be observed for all our approaches on the Inspec dataset, where the keyword labelling regime is very different from the labelling regime of the KP20k training set, on which the models were trained.
KP20k Inspec Krapivin KPTimes JPTimes openKP
F1@5 F1@10 F1@5 F1@10 F1@5 F1@10 F1@5 F1@10 F1@5 F1@10 F1@1 F1@3 F1@5
Unsupervised baselines
TF-IDF 7.2 9.4 16.0 24.4 6.7 9.3 17.9 15.1 26.6 22.9 19.6 22.3 19.6
TextRank 18.0 15.0 28.6 33.9 18.5 16.0 2.2 3.0 1.2 2.6 5.4 7.6 7.9
Large Language Model baselines
ChatGLM2-6b / / 25.1 30.1 / / / / / / 16.0 11.0 8.6
ChatGPT 23.2 10.8 35.2 33.9 12.8 11.7 27.9 / / / 20.8 20.4 16.6
GPT4o-mini 22.7 19.8 34.4 42.1 22.1 20.5 19.5 16.6 29.3 25.2 34.0 26.8 26.8
Gemini-1.5-flash-8B 26.5 23.6 35.9 42.1 26.6 25.9 20.4 26.9 31.8 27.5 33.1 25.5 25.6
Supervised baselines
CopyRNN 31.7 27.3 24.4 28.9 30.5 26.6 40.6 39.3 25.6 24.6 21.7 23.7 21.0
CatSeqD 34.8 29.8 27.6 33.3 32.5 28.5 42.4 42.4 23.8 23.8 / / /
GPT-2 27.5 27.8 41.3 46.9 25.3 25.3 42.1 42.3 33.1 33.6 / / /
GPT-2+BiLSTM-CRF 35.5 36.0 46.2 52.4 28.7 28.8 47.8 47.9 38.6 38.9 / / /
TNT-KID 33.6 33.8 46.0 53.6 31.0 32.0 48.5 48.5 35.9 36.1 / / /
JointKPE 41.7 34.4 35.2 35.0 36.0 29.2 / / / / 36.4 39.1 33.8
Diff-KPE 41.7 34.3 32.3 35.0 35.0 31.4 / / / / 37.8 38.5 32.7
HybridMatch / / / / / / / / / / 37.3 39.4 34.1
Our approaches - just pre-training
DeBERTa 35.2 35.3 14.5 14.5 28.8 28.9 55.6 55.7 37.4 37.5 34.1 42.9 43.2
DeBERTa MoE 33.9 34.0 14.7 14.8 28.5 28.8 56.4 56.5 39.1 39.3 34.7 42.7 43.0
DeBERTa RNN 33.6 33.7 13.9 13.9 27.5 27.7 53.2 53.2 36.8 36.9 33.4 39.7 39.8
DeBERTa MoE RNN 36.3 36.5 15.1 15.1 29.6 30.3 59.7 59.8 39.3 39.5 29.6 41.5 42.2
Table 5: Performance comparison between our approaches trained solely on KP20k, KPTimes, and OpenKP, and the baseline approaches. Only one seed is used in this scenario. Bold represents the best score according to a specific criterion on a specific dataset.